OpenStudy (anonymous):

Solve 3x – 5y = 15 for y.

OpenStudy (anonymous):

3x – 5y = 15 subtract 3x from both sides what you get???

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

how???

OpenStudy (anonymous):

i mean 4

OpenStudy (anonymous):

3x – 5y = 15 subtract 3x from both sides 3x-3x-5y = 15-3x -5y = 15-3x

OpenStudy (anonymous):

can you do the last step now??

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

then do it

OpenStudy (anonymous):

hint: divide both sides by -5 and you are done

OpenStudy (anonymous):

k